\begin{thebibliography}{99}

%-----------------------------( type Ia supernovae

%\cite{supernova Ia}
\bibitem{supernova Ia}
%
%\cite{Perlmutter:1997zf}
%\bibitem{Perlmutter:1997zf} 
  S.~Perlmutter {\it et al.}  %[Supernova Cosmology Project Collaboration],
  %``Discovery of a supernova explosion at half the age of the Universe and its cosmological implications,''
  Nature {\bf 391}, 51 (1998)
  [astro-ph/9712212].
%
%\cite{Riess:1998cb}
%\bibitem{Riess:1998cb} 
  A.~G.~Riess {\it et al.}  %[Supernova Search Team Collaboration],
  %``Observational evidence from supernovae for an accelerating universe and a cosmological constant,''
  Astron.\ J.\  {\bf 116}, 1009 (1998)
  [astro-ph/9805201].
%
%\cite{Perlmutter:1998np}
%\bibitem{Perlmutter:1998np} 
  S.~Perlmutter {\it et al.}  %[Supernova Cosmology Project Collaboration],
  %``Measurements of Omega and Lambda from 42 high redshift supernovae,''
  Astrophys.\ J.\  {\bf 517}, 565 (1999)
  [astro-ph/9812133].

%-----------------------------)



%-----------------------------( cosmic microwave background experiments

%\cite{cmb}
\bibitem{cmb}
%
%\cite{de Bernardis:2000gy}
%\bibitem{de Bernardis:2000gy} 
  P.~de Bernardis {\it et al.}  %[Boomerang Collaboration],
  %``A Flat universe from high resolution maps of the cosmic microwave background radiation,''
  Nature {\bf 404}, 955 (2000)
  [astro-ph/0004404].
%
%\cite{Hanany:2000qf}
%\bibitem{Hanany:2000qf} 
  S.~Hanany, P.~Ade, A.~Balbi, J.~Bock, J.~Borrill, A.~Boscaleri, P.~de Bernardis and P.~G.~Ferreira {\it et al.},
  %``MAXIMA-1: A Measurement of the cosmic microwave background anisotropy on angular scales of 10 arcminutes to 5 degrees,''
  Astrophys.\ J.\  {\bf 545}, L5 (2000)
  [astro-ph/0005123].
%
%\cite{Bennett:2003bz}
%\bibitem{Bennett:2003bz} 
  C.~L.~Bennett {\it et al.}  %[WMAP Collaboration],
  %``First year Wilkinson Microwave Anisotropy Probe (WMAP) observations: Preliminary maps and basic results,''
  Astrophys.\ J.\ Suppl.\  {\bf 148}, 1 (2003)
  [astro-ph/0302207].
%
%\cite{Spergel:2003cb}
%\bibitem{Spergel:2003cb} 
  D.~N.~Spergel {\it et al.}  %[WMAP Collaboration],
  %``First year Wilkinson Microwave Anisotropy Probe (WMAP) observations: Determination of cosmological parameters,''
  Astrophys.\ J.\ Suppl.\  {\bf 148}, 175 (2003)
  [astro-ph/0302209].

%-----------------------------)



%-----------------------------( large scale structure

%\cite{lss}
\bibitem{lss}
%
%\cite{Tegmark:2003ud}
%\bibitem{Tegmark:2003ud} 
  M.~Tegmark {\it et al.}  %[SDSS Collaboration],
  %``Cosmological parameters from SDSS and WMAP,''
  Phys.\ Rev.\ D {\bf 69}, 103501 (2004)
  [astro-ph/0310723].
%
%\cite{Seljak:2004xh}
%\bibitem{Seljak:2004xh} 
  U.~Seljak {\it et al.}  %[SDSS Collaboration],
  %``Cosmological parameter analysis including SDSS Ly-alpha forest and galaxy bias: Constraints on the primordial spectrum of fluctuations, neutrino mass, and dark energy,''
  Phys.\ Rev.\ D {\bf 71}, 103515 (2005)
  [astro-ph/0407372].

%-----------------------------)



%-----------------------------( baryon acoustic oscillations

%\cite{bao}
\bibitem{bao}
%
%\cite{Eisenstein:2005su}
%\bibitem{Eisenstein:2005su} 
  D.~J.~Eisenstein {\it et al.}  %[SDSS Collaboration],
  %``Detection of the baryon acoustic peak in the large-scale correlation function of SDSS luminous red galaxies,''
  Astrophys.\ J.\  {\bf 633}, 560 (2005)
  [astro-ph/0501171].
%
%\cite{Percival:2009xn}
%\bibitem{Percival:2009xn} 
  W.~J.~Percival {\it et al.}  %[SDSS Collaboration],
  %``Baryon Acoustic Oscillations in the Sloan Digital Sky Survey Data Release 7 Galaxy Sample,''
  Mon.\ Not.\ Roy.\ Astron.\ Soc.\  {\bf 401}, 2148 (2010)
  [arXiv:0907.1660 [astro-ph.CO]].

%-----------------------------)



%-----------------------------( dark matter and dark energy

%\cite{dark energy}
\bibitem{dark energy}
%
%\cite{Frieman:2008sn}
%\bibitem{Frieman:2008sn} 
  J.~Frieman, M.~Turner and D.~Huterer,
  %``Dark Energy and the Accelerating Universe,''
  Ann.\ Rev.\ Astron.\ Astrophys.\  {\bf 46}, 385 (2008)
  [arXiv:0803.0982 [astro-ph]].
%
%\cite{Caldwell:2009ix}
%\bibitem{Caldwell:2009ix} 
  R.~R.~Caldwell and M.~Kamionkowski,
  %``The Physics of Cosmic Acceleration,''
  Ann.\ Rev.\ Nucl.\ Part.\ Sci.\  {\bf 59}, 397 (2009)
  [arXiv:0903.0866 [astro-ph.CO]].
%
%\cite{Li:2011sd}
%\bibitem{Li:2011sd} 
  M.~Li, X.~-D.~Li, S.~Wang and Y.~Wang,
  %``Dark Energy,''
  Commun.\ Theor.\ Phys.\  {\bf 56}, 525 (2011)
  [arXiv:1103.5870 [astro-ph.CO]].

%-----------------------------)



%-----------------------------( f(R) first proposed

%\cite{f(R) first proposed}
\bibitem{f(R) first proposed}
%
%\cite{Buchdahl:1983zz}
%\bibitem{Buchdahl:1983zz} 
  H.~A.~Buchdahl,
  %``Non-linear Lagrangians and cosmological theory,''
  Mon.\ Not.\ Roy.\ Astron.\ Soc.\  {\bf 150}, 1 (1970).
%
%\cite{Kerner:1982yg}
%\bibitem{Kerner:1982yg} 
  RKerner,
  %``Cosmology without singularity and nonlinear gravitational Lagrangians,''
  Gen.\ Rel.\ Grav.\  {\bf 14}, 453 (1982).
%
%\cite{Barrow:1983rx}
%\bibitem{Barrow:1983rx} 
  J.~D.~Barrow and A.~C.~Ottewill,
  %``The Stability of General Relativistic Cosmological Theory,''
  J.\ Phys.\ A {\bf 16}, 2757 (1983).

%-----------------------------)



%-----------------------------( f(R) now

%\cite{f(R) now}
\bibitem{f(R) now}
%
%\cite{Nojiri:2006ri}
%\bibitem{Nojiri:2006ri} 
  S.~'i.~Nojiri and S.~D.~Odintsov,
  %``Introduction to modified gravity and gravitational alternative for dark energy,''
  eConf C {\bf 0602061}, 06 (2006)
  [Int.\ J.\ Geom.\ Meth.\ Mod.\ Phys.\  {\bf 4}, 115 (2007)]
  [hep-th/0601213].
%
%\cite{DeFelice:2010aj}
%\bibitem{DeFelice:2010aj} 
  A.~De Felice and S.~Tsujikawa,
  %``f(R) theories,''
  Living Rev.\ Rel.\  {\bf 13}, 3 (2010)
  [arXiv:1002.4928 [gr-qc]].
%
%\cite{Capozziello:2011et}
%\bibitem{Capozziello:2011et} 
  S.~Capozziello and M.~De Laurentis,
  %``Extended Theories of Gravity,''
  Phys.\ Rept.\  {\bf 509}, 167 (2011)
  [arXiv:1108.6266 [gr-qc]].
%
%\cite{Clifton:2011jh}
%\bibitem{Clifton:2011jh} 
  T.~Clifton, P.~G.~Ferreira, A.~Padilla and C.~Skordis,
  %``Modified Gravity and Cosmology,''
  Phys.\ Rept.\  {\bf 513}, 1 (2012)
  [arXiv:1106.2476 [astro-ph.CO]].
%
%\cite{Koivisto:2005yk}
%\bibitem{Koivisto:2005yk} 
  T.~Koivisto,
  %``Covariant conservation of energy momentum in modified gravities,''
  Class.\ Quant.\ Grav.\  {\bf 23}, 4289 (2006)
  [gr-qc/0505128].

%-----------------------------)



%-----------------------------( Severe weak field constraints

%\cite{wf}
\bibitem{wf}
%
%\cite{Nojiri:2003ft}
%\bibitem{Nojiri:2003ft} 
  S.~'i.~Nojiri and S.~D.~Odintsov,
  %``Modified gravity with negative and positive powers of the curvature: Unification of the inflation and of the cosmic acceleration,''
  Phys.\ Rev.\ D {\bf 68}, 123512 (2003)
  [hep-th/0307288].
%
%\cite{Cembranos:2005fi}
%\bibitem{Cembranos:2005fi} 
  J.~A.~R.~Cembranos,
  %``The Newtonian limit at intermediate energies,''
  Phys.\ Rev.\ D {\bf 73}, 064029 (2006)
  [gr-qc/0507039].
%
%\cite{Faulkner:2006ub}
%\bibitem{Faulkner:2006ub} 
  T.~Faulkner, M.~Tegmark, E.~F.~Bunn and Y.~Mao,
  %``Constraining f(R) Gravity as a Scalar Tensor Theory,''
  Phys.\ Rev.\ D {\bf 76}, 063505 (2007)
  [astro-ph/0612569].
%
%\cite{Hu:2007nk}
%\bibitem{Hu:2007nk} 
  W.~Hu and I.~Sawicki,
  %``Models of f(R) Cosmic Acceleration that Evade Solar-System Tests,''
  Phys.\ Rev.\ D {\bf 76}, 064004 (2007)
  [arXiv:0705.1158 [astro-ph]].
%
%\cite{Amendola:2007nt}
%\bibitem{Amendola:2007nt} 
  L.~Amendola and S.~Tsujikawa,
  %``Phantom crossing, equation-of-state singularities, and local gravity constraints in f(R) models,''
  Phys.\ Lett.\ B {\bf 660}, 125 (2008)
  [arXiv:0705.0396 [astro-ph]].

%-----------------------------)



%-----------------------------( f(R) gravity review

%\cite{f(R) review}
\bibitem{f(R) review}
%
%\cite{Sotiriou:2008rp}
%\bibitem{Sotiriou:2008rp} 
  T.~P.~Sotiriou and V.~Faraoni,
  %``f(R) Theories Of Gravity,''
  Rev.\ Mod.\ Phys.\  {\bf 82}, 451 (2010)
  [arXiv:0805.1726 [gr-qc]].
%
%\cite{DeFelice:2010aj}
%\bibitem{DeFelice:2010aj} 
  A.~De Felice and S.~Tsujikawa,
  %``f(R) theories,''
  Living Rev.\ Rel.\  {\bf 13}, 3 (2010)
  [arXiv:1002.4928 [gr-qc]].

%-----------------------------)



%-----------------------------( f(R) to f(R,L_m) step 1

%\cite{Nojiri:2004bi}
\bibitem{Nojiri:2004bi} 
  S.~'i.~Nojiri and S.~D.~Odintsov,
  %``Gravity assisted dark energy dominance and cosmic acceleration,''
  Phys.\ Lett.\ B {\bf 599}, 137 (2004)
  [astro-ph/0403622].
%
%\cite{Allemandi:2005qs}
%\bibitem{Allemandi:2005qs} 
  G.~Allemandi, A.~Borowiec, M.~Francaviglia and S.~D.~Odintsov,
  %``Dark energy dominance and cosmic acceleration in first order formalism,''
  Phys.\ Rev.\ D {\bf 72}, 063505 (2005)
  [gr-qc/0504057].

%\cite{study1}
\bibitem{study1}
%
%\cite{Bertolami:2007gv}
%\bibitem{Bertolami:2007gv} 
  O.~Bertolami, C.~G.~Bohmer, T.~Harko and F.~S.~N.~Lobo,
  %``Extra force in f(R) modified theories of gravity,''
  Phys.\ Rev.\ D {\bf 75}, 104016 (2007)
  [arXiv:0704.1733 [gr-qc]].
%
%\cite{Bertolami:2007vu}
%\bibitem{Bertolami:2007vu} 
  O.~Bertolami and J.~Paramos,
  %``Do f(R) theories matter?,''
  Phys.\ Rev.\ D {\bf 77}, 084018 (2008)
  [arXiv:0709.3988 [astro-ph]].
%
%\cite{Faraoni:2007sn}
%\bibitem{Faraoni:2007sn} 
  V.~Faraoni,
  %``A Viability criterion for modified gravity with an extra force,''
  Phys.\ Rev.\ D {\bf 76}, 127501 (2007)
  [arXiv:0710.1291 [gr-qc]].
%
%\cite{Sotiriou:2008dh}
%\bibitem{Sotiriou:2008dh} 
  T.~P.~Sotiriou,
  %``The Viability of theories with matter coupled to the Ricci scalar,''
  Phys.\ Lett.\ B {\bf 664}, 225 (2008)
  [arXiv:0805.1160 [gr-qc]].
%
%\cite{Bertolami:2008im}
%\bibitem{Bertolami:2008im} 
  O.~Bertolami and J.~Paramos,
  %``On the non-trivial gravitational coupling to matter,''
  Class.\ Quant.\ Grav.\  {\bf 25}, 245017 (2008)
  [arXiv:0805.1241 [gr-qc]].
%
%\cite{Puetzfeld:2008xu}
%\bibitem{Puetzfeld:2008xu} 
  D.~Puetzfeld and Y.~N.~Obukhov,
  %``On the motion of test bodies in theories with non-minimal coupling,''
  Phys.\ Rev.\ D {\bf 78}, 121501 (2008)
  [arXiv:0811.0913 [astro-ph]].
%
%\cite{Bertolami:2008zh}
%\bibitem{Bertolami:2008zh} 
  O.~Bertolami, J.~Paramos, T.~Harko and F.~S.~N.~Lobo,
  %``Non-minimal curvature-matter couplings in modified gravity,''
  arXiv:0811.2876 [gr-qc].
%
%\cite{Sotiriou:2008it}
%\bibitem{Sotiriou:2008it} 
  T.~P.~Sotiriou and V.~Faraoni,
  %``Modified gravity with R-matter couplings and (non-)geodesic motion,''
  Class.\ Quant.\ Grav.\  {\bf 25}, 205002 (2008)
  [arXiv:0805.1249 [gr-qc]].
%
%\cite{Bertolami:2008ab}
%\bibitem{Bertolami:2008ab} 
  O.~Bertolami, F.~S.~N.~Lobo and J.~Paramos,
  %``Non-minimum coupling of perfect fluids to curvature,''
  Phys.\ Rev.\ D {\bf 78}, 064036 (2008)
  [arXiv:0806.4434 [gr-qc]].
%
%\cite{Faraoni:2009rk}
%\bibitem{Faraoni:2009rk} 
  V.~Faraoni,
  %``The Lagrangian description of perfect fluids and modified gravity with an extra force,''
  Phys.\ Rev.\ D {\bf 80}, 124040 (2009)
  [arXiv:0912.1249 [astro-ph.GA]].
%
%\cite{Bertolami:2009cd}
%\bibitem{Bertolami:2009cd} 
  O.~Bertolami and M.~C.~Sequeira,
  %``Energy Conditions and Stability in f(R) theories of gravity with non-minimal coupling to matter,''
  Phys.\ Rev.\ D {\bf 79}, 104010 (2009)
  [arXiv:0903.4540 [gr-qc]].
%
%\cite{Nesseris:2008mq}
%\bibitem{Nesseris:2008mq} 
  S.~Nesseris,
  %``Matter density perturbations in modified gravity models with arbitrary coupling between matter and geometry,''
  Phys.\ Rev.\ D {\bf 79}, 044015 (2009)
  [arXiv:0811.4292 [astro-ph]].

%-----------------------------)



%-----------------------------( f(R) to f(R,L_m) step 2

%\cite{Harko:2010mv}
\bibitem{Harko:2010mv} 
  T.~Harko and F.~S.~N.~Lobo,
  %``f(R,$L_{m}$) gravity,''
  Eur.\ Phys.\ J.\ C {\bf 70}, 373 (2010)
  [arXiv:1008.4193 [gr-qc]].

%\cite{study2}
\bibitem{study2}
%
%\cite{Harko:2012ar}
%\bibitem{Harko:2012ar} 
  T.~Harko and F.~S.~N.~Lobo,
  %``Generalized dark gravity,''
  Int.\ J.\ Mod.\ Phys.\ D {\bf 21}, 1242019 (2012)
  [arXiv:1205.3284 [gr-qc]].
%
%\cite{Harko:2012hm}
%\bibitem{Harko:2012hm} 
  T.~Harko, F.~S.~N.~Lobo and O.~Minazzoli,
  %``Extended $f(R,L_m)$ gravity with generalized scalar field and kinetic term dependences,''
  Phys.\ Rev.\ D {\bf 87}, 047501 (2013)
  [arXiv:1210.4218 [gr-qc]].
%
%\cite{Harko:2012ve}
%\bibitem{Harko:2012ve} 
  T.~Harko and F.~S.~N.~Lobo,
  %``Geodesic deviation, Raychaudhuri equation, and tidal forces in modified gravity with an arbitrary curvature-matter coupling,''
  Phys.\ Rev.\ D {\bf 86}, 124034 (2012)
  [arXiv:1210.8044 [gr-qc]].
%
%\cite{Wang:2012tm}
%\bibitem{Wang:2012tm} 
  J.~Wang and K.~Liao,
  %``Energy conditions in $f(R,L_{m})$ gravity,''
  Class.\ Quant.\ Grav.\  {\bf 29}, 215016 (2012)
  [arXiv:1212.4656 [physics.gen-ph]].
%
%\cite{Harko:2013yb}
%\bibitem{Harko:2013yb} 
  T.~Harko, F.~S.~N.~Lobo, M.~K.~Mak and S.~V.~Sushkov,
  %``Gravitationally modified wormholes without exotic matter,''
  Phys.\  Rev.\ D {\bf 87}, 067504 (2013)
  [arXiv:1301.6878 [gr-qc]].

%-----------------------------)



%\cite{DeWitt:1967yk}
\bibitem{DeWitt:1967yk} 
  B.~S.~DeWitt,
  %``Quantum Theory of Gravity. 1. The Canonical Theory,''
  Phys.\ Rev.\  {\bf 160}, 1113 (1967).



%-----------------------------( quantum cosmology

%\cite{quancos}
\bibitem{quancos}
%
%\cite{Atkatz:1994hy}
%\bibitem{Atkatz:1994hy} 
  D.~Atkatz,
  %``Quantum cosmology for pedestrians,''
  Am.\ J.\ Phys.\  {\bf 62}, 619 (1994).
%
%\cite{Halliwell:1988nw}
%\bibitem{Halliwell:1988nw} 
  J.~J.~Halliwell,
  %``A Bibliography Of Papers On Quantum Cosmology,''
  Int.\ J.\ Mod.\ Phys.\ A {\bf 5}, 2473 (1990).
%
%\cite{Coule:2004qf}
%\bibitem{Coule:2004qf} 
  D.~H.~Coule,
  %``Quantum cosmological models,''
  Class.\ Quant.\ Grav.\  {\bf 22}, R125 (2005)
  [gr-qc/0412026].
%
%\cite{Kiefer:2006it}
%\bibitem{Kiefer:2006it} 
  C.~Kiefer,
  %``Quantum cosmology: Expectations and results,''
  Annalen Phys.\  {\bf 15}, 316 (2006).

%-----------------------------)



%-----------------------------( Wheeler's superspace

%\cite{Wheelersuper}
\bibitem{Wheelersuper}
%\cite{Wheeler:1968kg}
%\bibitem{Wheeler:1968kg}
  J.~A.~Wheeler,
  \emph{Einstein's Vision},
  Springer-Verlag, Berlin (1968).
%
%\cite{Ryan:1972sp}
%\bibitem{Ryan:1972sp}
  M.~Ryan,
  \emph{Hamiltonian Cosmology},
  Springer-Verlag, New York, (1972).
%
%\cite{Giulini:2009np}
%\bibitem{Giulini:2009np} 
  D.~Giulini,
  %``The Superspace of Geometrodynamics,''
  Gen.\ Rel.\ Grav.\  {\bf 41}, 785 (2009)
  [arXiv:0902.3923 [gr-qc]].

%-----------------------------)



%\cite{Vilenkin:1985md}
\bibitem{Vilenkin:1985md} 
  A.~Vilenkin,
  %``Classical And Quantum Cosmology Of The Starobinsky Inflationary Model,''
  Phys.\ Rev.\ D {\bf 32}, 2511 (1985).



%\cite{Misner:1974qy}
\bibitem{Misner:1974qy} 
  C.~W.~Misner, K.~S.~Thorne and J.~A.~Wheeler,
  \emph{Gravitation},
  San Francisco (1973).



%\cite{Harko:2010zi}
\bibitem{Harko:2010zi} 
  T.~Harko,
  %``The matter Lagrangian and the energy-momentum tensor in modified gravity with non-minimal coupling between matter and geometry,''
  Phys.\ Rev.\ D {\bf 81}, 044021 (2010)
  [arXiv:1001.5349 [gr-qc]].



%-----------------------------( Ostrogradsky scheme

%\cite{Ostrogradsky}
\bibitem{Ostrogradsky}
%
%\cite{Ostrogradsky:1850}
%\bibitem{Ostrogradsky:1850}
  M.~V.~Ostrogradsky,
  Mem.\ Acad.\ St. Petersbourg {\bf 6}, 385 (1850).
%
%\cite{GrosseKnetter:1993td}
%\bibitem{GrosseKnetter:1993td} 
  C.~Grosse-Knetter,
  %``Effective Lagrangians with higher derivatives and equations of motion,''
  Phys.\ Rev.\ D {\bf 49}, 6709 (1994)
  [hep-ph/9306321].

%-----------------------------)

%\cite{Kiefer:2008sw}
\bibitem{Kiefer:2008sw} 
  C.~Kiefer and B.~Sandhoefer,
  %``Quantum Cosmology,''
  arXiv:0804.0672 [gr-qc].

\end{thebibliography}

\documentclass[prd,showpacs,twocolumn]{revtex4}

\usepackage{hyperref}
\usepackage{amssymb}
\usepackage{amsmath}

\begin{document}

\title{The derivation of the Wheeler-DeWitt equation of $f(R,L_m)$ gravity in minisuperspace}

\author{Ru-Nan Huang}
 \email{huangrn@shu.edu.cn}
 \affiliation{Department of Physics, Shanghai University, Shanghai, 200444, P.R.China}

\date{\today}

\begin{abstract}
In this paper we present the Wheeler-DeWitt equation of $f(R,L_m)$ gravity in flat FRW universe, which is the first step of the study of quantum $f(R,L_m)$ cosmology. In the minisuperspace spanned by the FRW scale factor $a$ and the Ricci scalar $R$, we explicitly examine and verify the equivalence of the reduced action. We then canonical quantize the $f(R,L_m)$ model and derive the corresponding Wheeler-DeWitt equation.
\end{abstract}

\pacs{
 04.50.Kd, 
 04.20.Fy, 
 04.60.-m  
}

\maketitle


\section{Introduction}

From cosmic observations, such as the observations of type Ia supernovae \cite{supernova Ia}, cosmic microwave background radiation \cite{cmb}, large scale structure \cite{lss} and baryon acoustic oscillations \cite{bao}, we know that our universe is currently accelerating. This acceleration is due to the mysterious presence of dark energy, which accounts for 70\% of the cosmic energy content \cite{dark energy}. $f(R)$ gravity, first proposed in \cite{f(R) first proposed}, where the standard Einstein-Hilbert action is replaced by an arbitrary function of the Ricci scalar $R$, now acts as a promising way to solve the dark energy problem \cite{f(R) now}. Viable models that meet the severe weak field criteria in the solar system regime have been given in \cite{wf}. One can see \cite{f(R) review} for a review of $f(R)$ gravity.

Recently, Nojiri et al. generalized the $f(R)$ gravity by including in the action an explicit coupling between geometry and matter, i.e., they wrote the action as \cite{Nojiri:2004bi}
\begin{equation*}
 S=\int d^4x\sqrt{-g}
 \left\{\frac{1}{\kappa^2}R+\left(\frac{R}{\mu^2}\right)^\alpha L_d\right\}
\end{equation*}
where $L_d$ is the matter lagrangian density. The cosmological implications of such non-minimal coupling were widely investigated in \cite{study1}. It is revealed that the non-minimal coupling will give rise to an extra force on massive particles, which is orthogonal to the four velocities, and consequently cause non-geodesic motions. Harko and Lobo further extended such models to the maximal extension of the Einstein-Hilbert action, i.e., they wrote the action as
\begin{equation*}
 S=\int d^4x\sqrt{-g} f(R,L_m)
\end{equation*}
where $f(R,L_m)$ is an arbitrary function of the Ricci scalar $R$ and the matter lagrangian density $L_m$ \cite{Harko:2010mv}. The properties of $f(R,L_m)$ gravity are extensively studied in \cite{study2}.


Quantum gravity, proposed by DeWitt in his pioneering paper \cite{DeWitt:1967yk} in 1967, is an effort to provide a quantum theory of the gravitational field and to cast light on the study of cosmology \cite{quancos}. In the canonical quantization approach, which starts with a split of spacetime into space and time, the wave function $\Psi(g_{ij})$ is defined on a superspace, which, according to Wheeler, is an infinite-dimensional manifold whose points correspond to metrics on a three-dimensional spatial foliation, and acts as an arena where the development of the geometries of general relativity takes place \cite{Wheelersuper}. Quantum geometrodynamics and loop quantum gravity are the popular examples of canonical quantum gravity.

Our main purpose in this paper is to present the Wheeler-DeWitt equation of $f(R,L_m)$ gravity in the flat FRW universe. The Wheeler-DeWitt equation plays the role of Schr\"odinger equation for the wave function of the universe, and is at the heart of every canonical approach of quantum cosmology. Since the three-geometry of the flat FRW universe is homogeneous and isotropic, there are no inhomogeneous fluctuations of the metric. Thus the infinite number of gravitational degrees of freedom are reduced to just two variables, the FRW scale factor $a$ and the Ricci scalar $R$, and the resulting configuration space is called minisuperspace. Adopting the method proposed by Vilenkin in \cite{Vilenkin:1985md}, we present the Wheeler-DeWitt equation of $f(R,L_m)$ gravity, which serves as the first step of the study of quantum $f(R,L_m)$ cosmology. 


The paper is organized as follows. In Sec.~\ref{sec:fgravity} we give a brief review of $f(R,L_m)$ gravity developed in~\cite{Harko:2010mv}. Sec.~\ref{sec:redaction} are devoted to the derivation of the reduced action of $f(R,L_m)$ gravity in minisuperspace. We first write the action of $f(R,L_m)$ gravity in flat FRW universe via a Lagrangian multiplier and integrate by parts to get a reduced action. We then apply the variation priciple to obtain the reduced field equations. We explicitly verify that these reduced equations are the same as the field equation in Sec.~\ref{sec:fgravity} under flat FRW metric, which vindicates the equivalence of the reduced action and the original one. In Sec.~\ref{sec:wdequation} we canonical quantize the $f(R,L_m)$ gravity model in minisuperspace and obtain the corresponding Wheeler-DeWitt equation. Sec.~\ref{sec:conclusion} contains a discussion and the conclusions. We use the natural unit system $8\pi G=c=\hbar=1$ throughout this paper.


\section{Brief review of $f(R,L_m)$ gravity}\label{sec:fgravity}

The maximal extension of the Einstein-Hilbert action for non-minimal coupling gravity is given by~\cite{Harko:2010mv}
\begin{equation}\label{eqn:oriaction}
 S=\int d^4x\sqrt{-g}f(R,L_m)
\end{equation}
where $f(R,L_m)$ is an arbitrary function of the Ricci scalar $R$ and the matter Lagrangian density $L_m$. The energy-momentum tensor is defined as usual
\begin{equation}\label{eqn:emtensor}
 T_{\mu\nu}=-\frac{2}{\sqrt{-g}}\frac{\delta(\sqrt{-g}L_m)}{\delta g^{\mu\nu}}
\end{equation}
Assuming that $L_m$ depends only on the metric $g_{\mu\nu}$ and not on their derivatives, (\ref{eqn:emtensor}) can be reduced into
\begin{equation}\label{eqn:redemtensor}
 T_{\mu\nu}=g_{\mu\nu}L_m-2\frac{\partial L_m}{\partial g^{\mu\nu}}
\end{equation}

Varying the action (\ref{eqn:oriaction}) with respect to the metric $g^{\mu\nu}$ and using the expression of $T_{\mu\nu}$ given by (\ref{eqn:redemtensor}), one obtains the field equations of the $f(R,L_m)$ model
\begin{multline}\label{eqn:orifieldeq}
 f_R R_{\mu\nu}+(g_{\mu\nu}\square-\nabla_\mu\nabla_\nu)f_R-\frac{1}{2}(f-f_{L_m}L_m)g_{\mu\nu} \\
 =\frac{1}{2}f_{L_m}T_{\mu\nu} 
\end{multline}
where $\square=g^{\mu\nu}\nabla_\mu\nabla_\nu$ and
\begin{equation}
 f_R \equiv \frac{\partial f(R,L_m)}{\partial R}, \quad
 f_{L_m} \equiv \frac{\partial f(R,L_m)}{\partial L_m}
\end{equation}
Choosing $f(R,L_m)=\frac{1}{2}R+L_m$, i.e., the Einstein-Hilbert action, (\ref{eqn:orifieldeq}) can be reduced to the Einstein equation, $R_{\mu\nu}-\frac{1}{2}g_{\mu\nu}R=T_{\mu\nu}$, just as expected.

\smallskip

Using the following mathematical identities~\cite{Misner:1974qy}
\begin{equation}
 (\square\nabla_\mu-\nabla_\mu\square)f_R=R_{\mu\nu}\nabla^\nu f_R
\end{equation}
and
\begin{equation}
 \nabla^\mu G_{\mu\nu}=\nabla^\mu(R_{\mu\nu}-\frac{1}{2}g_{\mu\nu}R)=0
\end{equation}
one can prove that
\begin{equation}\label{eqn:idcovfR}
 \nabla^\mu\left[f_R R_{\mu\nu}
 +(g_{\mu\nu}\square-\nabla_\mu\nabla_\nu)f_R
 -\frac{1}{2}g_{\mu\nu}f\right]=0
\end{equation}
Taking the covariant derivative of (\ref{eqn:orifieldeq}) and using the identity (\ref{eqn:idcovfR}), one obtains the equation for the divergence of the energy-momentum tensor $T_{\mu\nu}$
\begin{equation}
 \nabla^\mu T_{\mu\nu}=2(\nabla^\mu\ln f_{L_m})\frac{\partial L_m}{\partial g^{\mu\nu}}
\end{equation}
It is shown that in modified gravity with non-minimal coupling between matter and geometry, both the matter Lagrangian and the energy-momentum tensor are completely determined by the form of the coupling \cite{Harko:2010zi}. Once the matter Lagrangian density $L_m$ is given, one can choose appropriate forms of $f(R,L_m)$ to obtain conservative models which give $\nabla^\mu T_{\mu\nu}=0$.


\section{Reduced Lagrangian in minisuperspace}\label{sec:redaction}

We now consider the $f(R,L_m)$ gravity in flat FRW universe filled with matter and write the metric as
\begin{equation}\label{eqn:FRW}
 ds^2=-dt^2+a^2(t)d\vec{x}^2
\end{equation}
which is homogeneous and isotropic and thus drasticlly reduced the gravitational degrees of freedom. This spacetime geometry is consistent with the present cosmological observations, such as the cosmic microwave background radiation, which has deviations from perfect smoothness on the order of $10^{-5}$ \cite{cmb}.

The scalar curvature for the flat FRW metric (\ref{eqn:FRW}) is
\begin{equation}\label{eqn:Ricci}
 R=6(\frac{\ddot{a}}{a}+\frac{\dot{a}^2}{a^2})
\end{equation}
Plugging (\ref{eqn:FRW}) and (\ref{eqn:Ricci}) directly into the $f(R,L_m)$ gravity model (\ref{eqn:oriaction}) will yield us a higher-order derivative Lagrangian, that is, a Lagrangian that contains the second derivative of $a$. Such a Lagrangian can be canonical quantized by the Ostrogradsky scheme (see, for example, \cite{Ostrogradsky}).

However, we shall adopt the method proposed by Vilenkin \cite{Vilenkin:1985md} and use a Lagrangian multiplier $\lambda$ to rewrite the action (\ref{eqn:oriaction}) under the flat FRW metric (\ref{eqn:FRW}), where the expression of $R$ (\ref{eqn:Ricci}) is inserted as a constraint
\begin{equation}\label{eqn:mulaction}
 S=\int dt \left\{a^3f(R,L_m)
 -\lambda\left(R-6(\frac{\ddot{a}}{a}+\frac{\dot{a}^2}{a^2})\right)\right\}
\end{equation}
The action (\ref{eqn:mulaction}) can be regarded as a dynamical system described by two independent variables, the scale factor $a$ and the scalar curvature $R$. Varying (\ref{eqn:mulaction}) with respect to $R$ gives the value of $\lambda$
\begin{equation}\label{eqn:lambda}
 \lambda=a^3f_R
\end{equation}
We substitute (\ref{eqn:lambda}) back into (\ref{eqn:mulaction}) and integrate by parts to remove the second derivative of $a$, then obtain
\begin{multline}\label{eqn:redaction}
 S=\int dt \left[ a^3(f-f_RR)-6a\dot{a}^2f_R \right. \\
 \left. -6a^2\dot{a}\dot{R}f_{RR}-6a^2\dot{a}^2f_{RL_m}\frac{\partial L_m}{\partial a} \right]
\end{multline}
where we have assumed, as reviewd in Sec.~\ref{sec:fgravity}, that the matter Lagrangian density $L_m$ depends on the metric $g_{\mu\nu}$ and not on its derivatives. The reduced Lagrangian is
\begin{multline}\label{eqn:redlag}
 L(a,\dot{a},R,\dot{R})= a^3(f-f_RR)-6a\dot{a}^2f_R \\
 -6a^2\dot{a}\dot{R}f_{RR}-6a^2\dot{a}^2f_{RL_m}\frac{\partial L_m}{\partial a}
\end{multline}
where
\begin{equation}
 f_{RR}=\frac{\partial^2 f(R,L_m)}{\partial R^2}, \quad
 f_{RL_m}=\frac{\partial^2 f(R,L_m)}{\partial L_m\partial R} 
\end{equation}
Thus we get the reduced Lagrangian in which the FRW scale factor $a$ and the Ricci scalar $R$ play the role of independent dynamical variables.

It is straight forward to verify that the reduced action (\ref{eqn:redaction}) is equivalent to the original $f(R,L_m)$ gravity action (\ref{eqn:oriaction}) under the flat FRW metric (\ref{eqn:FRW}).


The variation of (\ref{eqn:redaction}) with respect to the scale factor $a$ gives
\begin{equation}\label{eqn:redfieldeq}
\begin{split}
 \frac{\delta S}{\delta a}
 &= \frac{\partial L}{\partial a}-\frac{d}{dt}\frac{\partial L}{\partial \dot{a}} \\
 &= 3fa^2 - 3f_Ra^2R + 6f_R\dot{a}^2 + 12f_Ra\ddot{a} \\
 & \quad + 12f_{RR}a\dot{a}\dot{R} + 6f_{RR}a^2\ddot{R} + 6f_{RRR}a^2\dot{R}^2 \\
 & \quad + f_{L_m}\frac{\partial L_m}{\partial a}a^3 - f_{RL_m}\frac{\partial L_m}{\partial a}a^3R \\
 & \quad + 18f_{RL_m}\frac{\partial L_m}{\partial a}a\dot{a}^2 
         + 12f_{RL_m}\frac{\partial L_m}{\partial a}a^2\ddot{a} \\
 & \quad + 6f_{RL_m}\frac{\partial^2 L_m}{\partial a^2}a^2\dot{a}^2
         + 12f_{RL_mR}\frac{\partial L_m}{\partial a}a^2\dot{a}\dot{R} \\
 & \quad + 6f_{RL_mL_m}\left(\frac{\partial L_m}{\partial a}\right)^2a^2\dot{a}^2
\end{split}
\end{equation}
Using the following mathematical identities
\begin{equation}
 \square f_R=\frac{1}{\sqrt{-g}}\partial_\mu(\sqrt{-g}g^{\mu\nu}\partial_\nu f_R)
\end{equation}
and
\begin{equation}
 \nabla_\mu\nabla_\nu f_R
 = \partial_\mu\partial_\nu f_R -\Gamma_{\mu\nu}^\lambda\partial_\lambda f_R
\end{equation}
one can directly verify that (\ref{eqn:redfieldeq}) implies the space-space component of the field equation (\ref{eqn:orifieldeq}) under the flat FRW metric (\ref{eqn:FRW}).

The variation of (\ref{eqn:redaction}) with respect to the Ricci scalar $R$ gives
\begin{equation}
\begin{split}
 \frac{\delta S}{\delta R}
 &= \frac{\partial L}{\partial R}-\frac{d}{dt}\frac{\partial L}{\partial \dot{R}} \\
 &= -a^3f_{RR}\left[R-6\left(\frac{\ddot{a}}{a}+\frac{\dot{a}^2}{a^2}\right)\right]
\end{split}
\end{equation}
which implies (\ref{eqn:Ricci}), the expression for Ricci scalar under the flat FRW metric (\ref{eqn:FRW}).

Thus, we explicitly verify that the reduced action (\ref{eqn:redaction}) is equivalent to the original $f(R,L_m)$ gravity model in such a way that the variations with respect to its dynamical variables $a$ and $R$ give the correct equations of motion as that of the $f(R,L_m)$ action (\ref{eqn:oriaction}) under the flat FRW metric (\ref{eqn:FRW}).


\section{Wheeler-DeWitt equation}\label{sec:wdequation}

Now we can canonical quantize the reduced $f(R,L_m)$ action (\ref{eqn:redaction}) in minisuperspace, and give a detailed derivation of the corresponding Wheeler-DeWitt equation
\begin{equation}
 \hat{H}\Psi=0
\end{equation}
which plays the role of Schr\"{o}dinger equation for the wave function of the universe in canonical quantum gravity.

We first introduce the following two new variables to diagonalize the derivative part of the reduced Lagrangian (\ref{eqn:redlag})
\begin{equation}\label{eqn:newvar}
 u \equiv a\sqrt{f_R}, \quad v \equiv \frac{1}{2}\ln f_R
\end{equation}
in terms of which the reduced Lagrangian (\ref{eqn:redlag}) reads
\begin{multline}
 L(u,\dot{u},v,\dot{v}) \\
 =\left(\frac{u}{\sqrt{f_R}}\right)^3
 \left[-6f_R\left(\frac{\dot{u}^2}{u^2}-\dot{v}^2\right)+(f-f_RR)\right]
\end{multline}
where $R$ should be expressed in terms of $v$ through (\ref{eqn:newvar}) once the exact form of $f(R,L_m)$ is given. The corresponding conjugate momenta are
\begin{subequations}\label{eqn:uvcamomenta}
\begin{align}
 P_u &= \frac{\partial L}{\partial \dot{u}} = -12\frac{u\dot{u}}{\sqrt{f_R}} \\
 P_v &= \frac{\partial L}{\partial \dot{v}} = 12\frac{u^3}{\sqrt{f_R}}\dot{v}
\end{align}
\end{subequations}
and the resulting Hamiltonian is
\begin{equation}\label{eqn:uvHam}
\begin{split}
 H &= \dot{u}P_u+\dot{v}P_v-L \\
 &= -\frac{1}{24}\frac{\sqrt{f_R}}{u}P_u^2 + \frac{1}{24}\frac{\sqrt{f_R}}{u^3}P_v^2 
 -\left(\frac{u}{\sqrt{f_R}}\right)^3(f-f_RR)
\end{split}
\end{equation}

The DeWitt metric of the minisuperspace spanned by $A,B=\{u,v\}$ is
\begin{equation}\label{eqn:demetric}
 G_{AB} = \left( \begin{array}{cc}
 \displaystyle -12\frac{u}{\sqrt{f_R}} & 0 \\
 0 & \displaystyle 12\frac{u^3}{\sqrt{f_R}}
 \end{array}\right)
\end{equation}
by which the Hamiltonian (\ref{eqn:uvHam}) can be written as
\begin{equation}\label{eqn:DeHam}
 H=\frac{1}{2}G^{AB}P_AP_B + V(u,v)
\end{equation}
where $G^{AB}$ is the inverse DeWitt metric and
\begin{equation}
 V(u,v)= -\left(\frac{u}{\sqrt{f_R}}\right)^3(f-f_RR)
\end{equation}
Since the Hamiltonian (\ref{eqn:DeHam}) is classically constrained to vanish because of the time-time component of the field equation (\ref{eqn:orifieldeq}), the corresponding operator $\hat{H}$ in the quantum theory must annihilate the wave function.


Making the following replacement in the Hamiltonian (\ref{eqn:DeHam})
\begin{equation}
 P_u \rightarrow -i\frac{\partial}{\partial u}, \quad
 P_v \rightarrow -i\frac{\partial}{\partial v}
\end{equation}
with the covariant Laplace-Beltrami ordering \cite{Kiefer:2008sw}
\begin{equation}
 G^{AB}P_AP_B \rightarrow -\nabla_{LB}^2 = -\frac{1}{\sqrt{-G}}\partial_A(\sqrt{-G}G^{AB}\partial_B)
\end{equation}
we obtain the Hamiltonian operator
\begin{multline}
 \hat{H} = \frac{1}{24}\frac{\sqrt{f_R}}{u^2}\frac{\partial}{\partial u}u\frac{\partial}{\partial u}
 -\frac{1}{24}\frac{\sqrt{f_R}}{u^3}\frac{\partial^2}{\partial v^2} \\
 -\left(\frac{u}{\sqrt{f_R}}\right)^3(f-f_RR)
\end{multline}
Imposing $\hat{H}\Psi=0$ gives the Wheeler-DeWitt equation
\begin{multline}\label{eqn:wdequation}
 \left[\frac{1}{24}\frac{\sqrt{f_R}}{u^2}\frac{\partial}{\partial u}u\frac{\partial}{\partial u}
 -\frac{1}{24}\frac{\sqrt{f_R}}{u^3}\frac{\partial^2}{\partial v^2}\right. \\
 \left.-\left(\frac{u}{\sqrt{f_R}}\right)^3(f-f_RR)\right]\Psi(u,v)=0
\end{multline}
where, as stated before, the Ricci scalar $R$ should be expressed in terms of $v$ through (\ref{eqn:newvar}) once the exact form of $f(R,L_m)$ is given.

The Wheeler-DeWitt equation (\ref{eqn:wdequation}), which can be regarded as an analogue of the Schr\"odinger equation in quantum $f(R,L_m)$ gravity, is a partial differential equation of the wave function $\Psi(u,v)$ in minisuperspace. As a result of quantized diffeomorphism-invariant theory, equation (\ref{eqn:wdequation}) is defined on the geometrical and matter degrees of freedom of the model, and not on spacetime points. It contains terms that diverge as the big bang (or the big crunch) is approached, $u=a\sqrt{f_R}\rightarrow 0$ as expected.


\section{Conclusion}\label{sec:conclusion}

In this paper we have derived the Wheeler-DeWitt equation of $f(R,L_m)$ gravity in flat FRW universe, where the three-geometry is homogeneous and isotropic. Since we are not dealing with inhomogeneous fluctuations of the metric, the infinite number of gravitational degrees of freedom are reduced to just two variables, the FRW scale factor $a$ and the Ricci scalar $R$. The reduced action of $f(R,L_m)$ gravity in the truncated minisuperspace can be given via a Lagrangian multiplier and then integration by parts. The equivalence of the reduced action and the original action are explicitly verified. Using the canonical quantization procedure, the Wheeler-DeWitt equation of $f(R,L_m)$ gravity is obtained. The Wheeler-DeWitt equation plays the role of Schr\"odinger equation for the wave function of the universe, and is at the heart of every canonical approach of quantum cosmology. Thus our work can serve as the first step of the study of quantum $f(R,L_m)$ cosmology. \medskip


\section{Acknowledgments}

I would like to thank David Atkatz for emailing me his paper on quantum cosmology \cite{quancos}, and Tomi Koivisto for email correspondence. This work was supported by the National Natural Science Foundation of China (NSFC) under No.\ 11247309, the Excellent Young Teachers Program of Universities in Shanghai (No.\ shu11028), and the Innovating Funds of Shanghai university.


\end{document}